Benitez Company currently outsources a relay switch that is a component in one of its products. The switches cost $23 each. The company is considering making the switches internally at the following projected annual production costs:

Unit-level material cost $ 4
Unit-level labor cost $ 3
Unit-level overhead $ 2
Batch-level set-up cost (6,000 units per batch) $ 32,000
Product-level supervisory salaries $ 41,000
Allocated facility-level costs $ 27,000

The company expects an annual need for 6,000 switches. If the company makes the product, it will have to utilize factory space currently being leased to another company for $2,200 a month. If the company decides to make the parts, total costs will be:

Multiple Choice

  • $25,600 less than if the switches are purchased.

  • $15,400 more than if the switches are purchased.

  • $42,400 more than if the switches are purchased.

  • $27,000 less than if the switches are purchased.
Answer #1

The correct answer is

$15400 more than if the switches are purchased

Explanation

Total cost of manufacture

= total variable cost + batch level cost + supervisor salary + loss of lease rent

= (4+3+2)×6000 + 32000+ 41000 +(2200×12)

= 153400

Total purchase cost = 23×6000 = 138000

Thus manufacturing relevant cost is more by $15400 (153400 - 138000)